The base of a triangle is x cm long and its altitude is quarter the length of its base. if the triangle has an area of 12 cm2, form an equation in x and solve it.
Answer by nerdybill(7384)

The base of a triangle is x cm long and its altitude is quarter the length of its base. if the triangle has an area of 12 cm2, form an equation in x and solve it.
.
altitude is the same as the height of a triangle
.
and, we know that for any triangle
area = (1/2)bh
where
b is base
h is height
.
Let x = base
then
x/4 = altitude
12 = (1/2)x(x/4)
12 = x(x/8)
12 = x^2/8
12/8 = x^2
3/2 = x^2
1.225 cm = x (base)
.
altitude:
.306 cm